College Admissions:

Hi, I am looking to get into an impacted program. The schools that offer it have, on average, 300 qualified applicants for 60 slots. All 300 qualifying applications go into a lottery.

If I manage to become one of these 300 qualified applicants, and I apply at 5 schools, all of which use the same process to distribute the slots (same availability of slots, same number of qualifying applications), what is my chance of getting into the program at one of these schools in percent?

Can someone show me how to work this out?

Many thanks in advance! You're work is highly appreciated!

Answer by edjones(7311)

60/300=1/5 chances of being selected by one program
4/5 chance of not being selected by 1 program.
(4/5)^5 chance of not being selected by any of the 5 programs.
(4/5)^5=16/25
1- 16/25=9/25=.36=36% chances of getting into the program in at least one of the 5 schools.
.
Ed